# Zero-Shot Statistical Tests for LLM-Generated Text Detection using Finite Sample Concentration Inequalities

### v3

@@ 6 Conclusion
In this study, we establish first zero-shot statistical tests with theoretical guarantees for text with finite length to distinguish between (i) LLM-generated and human-generated texts in Proposition 3, and (ii) the text generated by two different LLMs $A$ and $B$ in Propositions 1 and 2. We prove that the type I and type II errors for our tests decrease exponentially in the text length. As a critical step in designing our tests, we derive concentration bounds in the difference between log-perplexity and the average entropy of the string under $A$. Specifically, for a given string, in Theorem 1, we demonstrate that if the string is generated by $A$, the log-perplexity of the string under $A$ converges to the average entropy of the string under $A$, except with an exponentially small probability in string length. Furthermore, in Theorem 2, we show that if $B$ (which can be either another model or human) generates the text, then, except with an exponentially small probability in string length, the log-perplexity of the string under $A$ converges to the average cross-entropy of $B$ and $A$. Our theoretical results rely on establishing concentration bounds for the difference between the log-likelihood of a sequence of discrete random variables and the negative entropy for non-independent random variables on a finite alphabet. Results in the literature (e.g., Zhao (2022)) derive concentration bounds for iid random variables, and one of our theoretical contributions is to extend the results to non-independent random variables by introducing random variables that form a martingale. We hope that our work inspires more research on zero-shot LLM-text detection with provable guarantees.
@@ Appendix A Appendix
A.1 Proof of Lemma 2
Proof.
$\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})e^{\frac{\big{|}-\log p^{A}_{n}(y_{n})-\mathbb{E}[Z_{n}]\big{|}}{t}}\leq e^{\frac{\mathbb{E}[Z_{n}]}{t}}\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\frac{1}{p^{A}_{n}(y_{n})^{1/t}}=e^{\frac{\mathbb{E}[Z_{n}]}{t}}\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})^{\frac{t-1}{t}}\leq 2.$ (6)
To analyze $e^{\frac{\mathbb{E}[Z_{n}]}{t}}$, we first want to show that
$\mathbb{E}[Z_{n}]$ $=$ $-\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})$ (8)
$\leq$ $\log|\mathcal{X}|=\log(K).$
To show that, note that by concavity of $-p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})$, we can maximize $p_{n}(y_{n})\log p_{n}(y_{n})$ by equalizing all probabilities $p^{*}(y_{n})=\frac{1}{|\mathcal{X}|};\forall n$. So, we have
$=$ $-|\mathcal{X}|\Big{[}\frac{1}{|\mathcal{X}|}\log\frac{1}{|\mathcal{X}|}\Big{]}=\log|\mathcal{X}|.$
$K^{1/t}K^{1/t}\leq 2$
Hence, $X_{n}$ has sub-exponential norm $\|X\|_{\psi_{1}}=2\log(K)$. ∎
A.2 Proof of Theorem 1
Proof.
We first present an equivalent (up to a constant factor) definition of a sub-exponential random variable from Vershynin (2018).
@@ A.2 Proof of Theorem 1
Next, we present a lemma from Vershynin (2018) that demonstrates that the two definitions are equivalent up to a constant factor.
@@
(SE properties ,Vershynin, 2018). Let $X$ be a random variable with $\mathbb{E}[X]=0$. Then, there exists a constant $c$ and constants $K_{4}$ and $K_{5}$ such that $K_{4}\leq cK_{5}$ and $K_{5}\leq cK_{4}$ and the following two properties are equivalent.
@@
$\mathbb{E}[e^{\lambda X_{n}}|\mathbf{Y}_{n-1}]\leq e^{\frac{\nu^{2}\lambda^{2}}{2}}\quad\forall|\lambda|\leq\frac{1}{\nu}.$ (9)
Let $\alpha=\nu$. By Definition 2, we have $X_{n}\in SE(\nu^{2},\alpha)$.
Our next step is to show that $\sum_{n=1}^{N}X_{n}$ is SE with parameters $(\nu\sqrt{N},\alpha_{*})$. To realize that, observe that
$\mathbb{E}\left[e^{\lambda\left(\sum_{k=1}^{n}X_{k}\right)}\right]=\mathbb{E}\left[e^{\lambda\left(\sum_{k=1}^{n-1}X_{k}\right)}\mathbb{E}\left[e^{\lambda X_{n}}\mid\mathbf{Y}_{n-1}\right]\right]\leq\mathbb{E}\left[e^{\lambda\sum_{k=1}^{n-1}X_{k}}\right]e^{\frac{\lambda^{2}\nu^{2}}{2}}\leq e^{\frac{\lambda^{2}N\nu^{2}}{2}},$
Finally, from Theorem 5.2 in Arinaldo, 2019, if $S\in SE(\nu^{2},\alpha)$ is a sub-exponential random variable, then
Substituting $\nu=\alpha\sqrt{N}$ for the zero-mean random variable $\sum_{n=1}^{N}X_{n}$, and $\alpha=c_{1}\log K$, we obtain
$\mathbb{P}(|\sum_{n=1}^{N}X_{n}|\geq t_{1})\leq 2\exp\bigg{(}-\frac{1}{2}\min\bigg{(}\frac{t_{1}^{2}}{N(c_{1}\log K)^{2}},\frac{t_{1}}{c_{1}\log K}\bigg{)}\bigg{)}.$
Setting $t_{1}=tN$, we have
$\mathbb{P}\bigg{(}\frac{1}{N}|\sum_{i=1}^{N}X_{i}|\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nt}{c_{1}\log(K)}\min\bigg{(}1,\frac{t}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
A.3 Proof of Lemma 3
@@
$\sum_{y_{n}\in\mathcal{X}}p^{B}_{n}(y_{n})e^{\frac{\big{|}-\log p^{A}_{n}(y_{n})-\mathbb{E}[Z_{n}]\big{|}}{t}}\leq e^{\frac{\mathbb{E}[Z_{n}]}{t}}\sum_{y_{n}\in\mathcal{X}}\frac{p^{B}_{n}(y_{n})}{p^{A}_{n}(y_{n})^{1/t}}\leq e^{\frac{\mathbb{E}[Z_{n}]}{t}}\sum_{y_{n}\in\mathcal{X}}\frac{p^{B}_{n}(y_{n})}{\epsilon^{1/t}}\leq e^{\frac{\mathbb{E}[Z_{n}]-\log(\epsilon)}{t}}\leq 2.$
@@
$\frac{\mathbb{E}[Z_{n}]-\log(\epsilon)}{t}\leq\log(2)\leq 1/2.$ (11)
@@
$\mathbb{E}_{p^{B}_{n}}[Z_{n}]=\sum_{y_{n}\in\mathcal{X}}-p^{B}_{n}(y_{n})\log(p^{A}_{n}(y_{n}))\leq-\log(\epsilon).$
Substituting the result of (5) in (11) yields that $X_{n}$ has sub-exponential norm $\|X\|_{\psi_{1}}=-4\log(\epsilon)$. ∎
A.4 Proof of Theorem 2
@@
Following the same steps as in proof of Theorem 1, we conclude that $\sum_{n=1}^{N}X_{n}$ is sub-exponential with parameter $S_{N}\in SE(\nu\sqrt{N},\alpha)$, where $\alpha=\nu=-c_{3}\log(\epsilon)$ for a constant $c_{3}>0$. Then, from (10) we have
$\mathbb{P}(|\sum_{n=1}^{N}X_{n}|\geq t_{1})\leq 2\exp\bigg{(}-\frac{1}{2}\min\bigg{(}\frac{t_{1}^{2}}{N(-c_{3}\log\epsilon)^{2}},\frac{t_{1}}{-c_{3}\log\epsilon}\bigg{)}\bigg{)}.$
@@
$\mathbb{P}\bigg{(}\frac{1}{N}|\sum_{n=1}^{N}X_{n}|\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nt}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{t}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
A.5 Proof of Proposition 1

### v4

@@ 6 Conclusion
We answer the following question: Given a piece of text, can we identify whether it was produced by a particular LLM or not? We model LLM-generated text as a sequential stochastic process with complete dependence on history and design statistical tests that take a single finite length string of text, a prompt, and a given LLM as input and assess whether the given LLM produced the text. We first design composite tests that determine whether the text is generated by a model that belongs to a set of models $\mathcal{A}$ or to a disjoint set of models $\mathcal{B}$. We then design a composite statistical test to identify whether a text is generated by the model $A$ or not. In Propositions 1 and 2, we prove that our tests’ type I and type II errors decrease exponentially as the text length increases. As a prerequisite for designing our tests, in Theorem 1, we prove an exponential decay concentration inequality that bounds the tail probability of the difference between the log-perplexity of a sequence generated by $B$ and evaluated under $A$, and the average cross-entropy of $B$ and $A$, over a finite alphabet. The special case where (i) the sequence is independent and identically distributed (iid), and (ii) models $B$ and $A$ are the same, is studied in [72]. However, LLMs generate text sequences that depend on the previously generated tokens. Therefore, our result in Theorem 1 generalizes the results in [72], relaxing (i) and (ii). Interestingly, we address one of the future research directions mentioned by [72], which is considering a sequence of non-independent random variables. Our numerical experiments in a white-box setting show the validity of our theoretical results. In the black-box setting, our method achieves an average TPR of 82.5% at a fixed FPR of 5%. Under adversarial perturbations, the minimum TPR is 48.6% at the same FPR threshold. Both results outperform all non-commercial baselines. We hope our work inspires more research on zero-shot LLM-text detection with provable guarantees.
@@ Appendix A Appendix
A.1 Concentration inequality proofs
In this section, to improve readability, we first provide the prerequisites of sub-exponential random variables that are used in our proofs in Section A.1.1, followed by a review of notations in Section A.1.2. We provide the proof for Theorem 1 - part (a), i.e., where the generator model $B$ and the evaluator model $A$ are the same, in Section A.1.3. The starting paragraph in the proof provides a high-level overview, describing the proof steps, remarks, and lemmas used as part of the proof. In Section A.1.4, we provide the proof for Theorem 1 - part (b), i.e., where the generator model $B$ is not the same as the evaluator model $A$. Similar to the previous part, the starting paragraph of this section provides a proof overview.
A.1.1 Prerequisites of sub-exponential random variables from the literature
Before presenting the proofs for our main theorem, we need to clearly state the following definitions and the equivalency lemma for sub-exponential random variables from the literature.
Definition 1.
If $\|X\|_{\psi_{1}}$ is finite, we say that $X$ is sub-exponential.
We also need the following equivalent (up to a constant factor) definition of a sub-exponential random variable from [67].
@@ A.1.1 Prerequisites of sub-exponential random variables from the literature
Next, we present a lemma from Vershynin [67] that demonstrates that the two definitions are equivalent up to a constant factor.
@@
(SE properties ,67). Let $X$ be a random variable with $\mathbb{E}[X]=0$. Then, there exists a constant $c$ and constants $K_{4}$ and $K_{5}$ such that $K_{4}\leq cK_{5}$ and $K_{5}\leq cK_{4}$ and the following two properties are equivalent.
@@
A.1.2 Review of notations
Here, we first review the notation in Section 2: X denotes user prompt and Y denotes the output that consists of a string of tokens $\textbf{Y}=[Y_{1},Y_{2},\dots,Y_{N},\dots]$. Each token is chosen from a finite vocabulary set, i.e., $Y_{n}\in\mathcal{X}$ with size $K:=|\mathcal{X}|$. We defined the random variable $Z_{n}^{A}:=-\log\big{(}p^{A}_{n}(Y_{n})\big{)}$, and the zero-mean random variable $X_{n}^{A}:=Z_{n}^{A}-\mathbb{E}_{p^{B}_{n}}[Z_{n}^{A}]$. Additionally, we defined a random variable $S_{N}^{A}:=\sum_{i=1}^{N}X^{A}_{i}$.
A.1.3 Proof of Theorem 1 - part (a)
We first provide a high-level structure for our proof. Recall that in part (a), we focus on the case where the generator model $B$ is the same as the evaluator model $A$. In Remark 1, we show that $\mathbb{E}_{p_{n}^{A}}[Z_{n}^{A}]$ is upper bounded by $\log K$, and therefore it is finite. In Lemma 2, we show that the random variable $S_{N}^{A}=\sum_{i=1}^{N}X_{i}^{A}$ forms a martingale. In Lemma 3, we find the sub-exponential norm for the martingale increments random variable $X_{n}^{A}$. As the last step, we show that martingale sum $\sum_{n=1}^{N}X_{n}^{A}$ is sub-exponential with parameters $(\nu^{2}N,\nu)$, with $\nu=c_{1}\log K$. Finally, we apply the tail bound for sub-exponential random variables to bound the probability of $|\sum_{n=1}^{N}X_{n}^{A}|$ being larger than a constant $t$. A challenge in applying the common concentration bounds for martingales is that martingale increments are not necessarily bounded. We overcome this issue by showing that the martingale differences, while not bounded, admit a light tail. In particular, we show that the martingale differences are sub-exponential.
Remark 1.
$\mathbb{E}_{p_{n}^{A}}[Z_{n}^{A}]=-\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})\leq\log K.$
Proof.
Concavity of $-p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})$ yields that its maximizer is $p^{*A}_{n}(y_{n})=\frac{1}{K},\forall y_{n}\in\chi$. Hence,
$\mathbb{E}_{p_{n}^{A}}[Z_{n}^{A}]=-\sum_{y_{n}\in\chi}p_{n}^{A}(y_{n})\log p_{n}^{A}(y_{n})$$\leq$ $\sum_{y_{n}\in\chi}p_{n}^{*A}(y_{n})\log p_{n}^{*A}(y_{n})$
$=$ $K.\frac{1}{K}.\log K=\log K.$
Now that we have shown $\mathbb{E}_{p_{n}^{A}}[Z_{n}^{A}]<\infty$, recall that we defined the zero-mean random variable $X_{n}^{A}=Z_{n}^{A}-\mathbb{E}_{p_{n}^{A}}[Z_{n}^{A}]$. Then, we defined $S_{N}^{A}=\sum_{i=1}^{N}X_{i}^{A}$, which as we show in Lemma 2, forms a martingale.
Lemma 2.
If the evaluator model $A$ is the same as the generator model $B$ that generates the string $\bf{Y}_{N}$, then the random variable $S_{N}=\sum_{i=1}^{N}X_{i}^{A}$ forms a martingale.
Proof.
$Z_{n}^{A}$ and as a result $\mathbb{E}_{p_{n}^{A}}(y_{n})$ are positive random variables. In Remark 1, we showed that $\mathbb{E}_{p_{n}^{A}}[Z_{n}]\leq\log K$. Therefore, we have $\mathbb{E}\big{[}|X_{n}^{A}|\big{]}\leq\log K$. For a random variable to form a martingale, the following two properties need to be satisfied: (i) $\mathbb{E}\big{[}|S_{N+1}^{A}-S_{N}^{A}|\big{]}<\infty$, and (ii) $\mathbb{E}[S_{N+1}^{A}|S_{N}^{A}]=S_{N}^{A}$.
Property (i) is satisfied because $\mathbb{E}\big{[}|S_{N+1}^{A}-S_{N}^{A}|\big{]}=\mathbb{E}\big{[}|X_{N+1}^{A}|\big{]}\leq\log K<\infty$.
Property (ii) is satisfied because the martingale increments $X_{n}^{A}$ are, by definition, a zero-mean random variable conditioned on past tokens. ∎
Lemma 3.
The sub-exponential norm for the random variable $X_{n}^{A}$ equals $2\log(K)$.
@@ A.1.3 Proof of Theorem 1 - part (a)
$\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})e^{\frac{\big{|}-\log p^{A}_{n}(y_{n})-\mathbb{E}[Z_{n}^{A}]\big{|}}{t}}\leq e^{\frac{\mathbb{E}[Z_{n}]}{t}}\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\frac{1}{p^{A}_{n}(y_{n})^{1/t}}=e^{\frac{\mathbb{E}[Z_{n}]}{t}}\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})^{\frac{t-1}{t}}\leq 2.$ (6)
To analyze $e^{\frac{\mathbb{E}[Z_{n}^{A}]}{t}}$, we first want to show that
$\mathbb{E}[Z_{n}^{A}]$ $=$ $-\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})$ (8)
$\leq$ $\log(K).$
To show that, note that by concavity of $-p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})$, we can maximize $p_{n}(y_{n})\log p_{n}(y_{n})$ by equalizing all probabilities $p^{*}(y_{n})=\frac{1}{K};\forall n$. So, we have
$=$ $-K\Big{[}\frac{1}{K}\log\frac{1}{K}\Big{]}=\log K.$
$K^{1/t}K^{1/t}\leq 2$
Hence, $X_{n}^{A}$ has sub-exponential norm $\|X\|_{\psi_{1}}=2\log(K)$. ∎
Proof of Theorem 1 part (a).
Proof.
$\mathbb{E}[e^{\lambda X_{n}^{A}}|\mathbf{Y}_{n-1}]\leq e^{\frac{\nu^{2}\lambda^{2}}{2}}\quad\forall|\lambda|\leq\frac{1}{\nu}.$ (9)
By Definition 2, we have $X_{n}\in SE(\nu^{2},\nu)$.
Our next step is to show that $\sum_{n=1}^{N}X_{n}$ is SE with parameters $(\nu^{2}N,\nu)$. To realize that, observe that
$\mathbb{E}\left[e^{\lambda\left(\sum_{k=1}^{n}X_{k}^{A}\right)}\right]=\mathbb{E}\left[e^{\lambda\left(\sum_{k=1}^{n-1}X_{k}^{A}\right)}\mathbb{E}\left[e^{\lambda X_{n}^{A}}\mid\mathbf{Y}_{n-1}\right]\right]\leq\mathbb{E}\left[e^{\lambda\sum_{k=1}^{n-1}X_{k}^{A}}\right]e^{\frac{\lambda^{2}\nu^{2}}{2}}\leq e^{\frac{\lambda^{2}N\nu^{2}}{2}},$
Finally, from Theorem 5.2 in 2, if $S\in SE(\nu^{2},\alpha)$ is a sub-exponential random variable, then
Substituting for the zero-mean random variable $\sum_{n=1}^{N}X_{n}$ that is sub-exponential with parameters $(\nu^{2}N,\nu)$, with $\nu=c_{1}\log K$, we obtain
$\mathbb{P}(|\sum_{n=1}^{N}X_{n}^{A}|\geq t_{1})\leq 2\exp\bigg{(}-\frac{1}{2}\min\bigg{(}\frac{t_{1}^{2}}{N(c_{1}\log K)^{2}},\frac{t_{1}}{c_{1}\log K}\bigg{)}\bigg{)}.$
Setting $t_{1}=tN$, we have
$\mathbb{P}\bigg{(}\frac{1}{N}|\sum_{i=1}^{N}X_{i}^{A}|\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nt}{c_{1}\log(K)}\min\bigg{(}1,\frac{t}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
This concludes the proof for Theorem 1 part (a). ∎
A.1.4 Proof of Theorem 1 - part (b)
Similar to the proof of part (a), here we provide a high-level structure for our proof. First, recall from the discussion under Assumption 1, as proved in Equation 5, that $\mathbb{E}_{p_{n}^{A}}[Z_{n}^{A}]$ is upper bounded by $-\log(\epsilon)$, and therefore it is finite. Then, In Lemma 4, we show that the random variable $S_{N}^{A}=\sum_{i=1}^{N}X_{i}^{A}$ forms a martingale. In Lemma 5, we find the sub-exponential norm for the martingale increments random variable $X_{n}^{A}$. As the last step, we show that martingale sum $\sum_{n=1}^{N}X_{n}^{A}$ is sub exponential with parameters $(\nu^{2}N,\nu)$, where $\nu=-c_{3}\log(\epsilon)$. Finally, we apply the tail bound for sub-exponential random variables to bound the probability of $|\sum_{n=1}^{N}X_{n}^{A}|$ being larger than a constant $t$.
Lemma 4.
Under assumption 1, the random variable $S_{N}^{A}=\sum_{i=1}^{N}X_{i}^{A}$ forms a martingale.
Proof.
$Z_{n}^{A}$ is a positive random variables, so $\mathbb{E}_{p_{n}^{A}}[Z_{n}^{A}]$ is also positive. In Equation (5), we have shown that $\mathbb{E}_{p_{n}^{A}}[Z_{n}^{A}]$ is upper bounded by $-\log(\epsilon)$, and therefore we have $\mathbb{E}\big{[}|X_{n}^{A}|\big{]}\leq-\log(\epsilon)$. For a random variable to form a martingale, the following two properties need to be satisfied: (i) $\mathbb{E}\big{[}|S_{N+1}^{A}-S_{N}^{A}|\big{]}<\infty$, and (ii) $\mathbb{E}[S_{N+1}^{A}|S_{N}^{A}]=S_{N}^{A}$.
Property (i) is satisfied because $\mathbb{E}\big{[}|S_{N+1}^{A}-S_{N}^{A}|\big{]}=\mathbb{E}\big{[}|X_{N+1}^{A}|\big{]}\leq-\log(\epsilon)<\infty$.
Property (ii) is satisfied because the martingale increments $X_{n}^{A}$ are, by definition, a zero-mean random variable conditioned on past tokens. ∎
Lemma 5.
Proof.
$\sum_{y_{n}\in\mathcal{X}}p^{B}_{n}(y_{n})e^{\frac{\big{|}-\log p^{A}_{n}(y_{n})-\mathbb{E}[Z_{n}^{A}]\big{|}}{t}}\leq e^{\frac{\mathbb{E}[Z_{n}^{A}]}{t}}\sum_{y_{n}\in\mathcal{X}}\frac{p^{B}_{n}(y_{n})}{p^{A}_{n}(y_{n})^{1/t}}\leq e^{\frac{\mathbb{E}[Z_{n}^{A}]}{t}}\sum_{y_{n}\in\mathcal{X}}\frac{p^{B}_{n}(y_{n})}{\epsilon^{1/t}}\leq e^{\frac{\mathbb{E}[Z_{n}^{A}]-\log(\epsilon)}{t}}\leq 2.$
@@ A.1.4 Proof of Theorem 1 - part (b)
$\frac{\mathbb{E}[Z_{n}^{A}]-\log(\epsilon)}{t}\leq\log(2)\leq 1/2.$ (11)
@@
$\mathbb{E}_{p^{B}_{n}}[Z_{n}^{A}]=\sum_{y_{n}\in\mathcal{X}}-p^{B}_{n}(y_{n})\log(p^{A}_{n}(y_{n}))\leq-\log(\epsilon).$
Substituting the result of (5) in (11) yields that $X_{n}^{A}$ has sub-exponential norm $\|X\|_{\psi_{1}}=-4\log(\epsilon)$. ∎
Proof of Theorem 1 part (b).
@@
First note that, by Lemma 5 and Lemma 1, there exists $c_{3}>0$ s.t. for $\nu=-c_{3}\log(\epsilon)$, and for any $n\in\mathbb{N}$,
$\mathbb{E}[e^{\lambda X_{n}^{A}}|\mathbf{Y}_{n-1}]\leq e^{\frac{\nu^{2}\lambda^{2}}{2}}\quad\forall|\lambda|\leq\frac{1}{\nu}.$ (12)
By Definition 2, we have $X_{n}\in SE(\nu^{2},\nu)$. Then, Following the exact same steps as in proof of Theorem 1-part (a), we conclude that $\sum_{n=1}^{N}X_{n}^{A}$ is sub-exponential with parameter $S_{N}^{A}\in SE(\nu^{2}N,\nu)$, where $\nu=-c_{3}\log(\epsilon)$ for a constant $c_{3}>0$. Again, following the exact same steps we described in proof of part (a) , from (10) we have
$\mathbb{P}(|\sum_{n=1}^{N}X_{n}^{A}|\geq t_{1})\leq 2\exp\bigg{(}-\frac{1}{2}\min\bigg{(}\frac{t_{1}^{2}}{N(-c_{3}\log\epsilon)^{2}},\frac{t_{1}}{-c_{3}\log\epsilon}\bigg{)}\bigg{)}.$
@@
$\mathbb{P}\bigg{(}\frac{1}{N}|\sum_{n=1}^{N}X_{n}^{A}|\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nt}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{t}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
This concludes the proof for Theorem 1 part (b). ∎
A.2 Statistical test proofs
In Section A.2.1, we present the proof for Proposition 1. In Section A.2.2, we first provide a proof overview in the first paragraph, and then provide the proof for Proposition 2.
A.2.1 Proof of Proposition 1